The rms speed of the molecules of a gas at 147°C is 191 m/s. Calculate the mass of a single molecule in the gas.
Answer #1

Given

Vrms = 191 m/s

T = 147 C = 147 + 273 = 420 k

Let M is the molar mass of the gas.

we know, Vrms = sqrt(3*R*T/M)

Vrms^2 = 3*R*T/M

==> M = 3*R*T/Vrms^2

= 3*8.324*420/191^2

= 0.2875 kg/mol

mass of one moelcule, m = M/NA

= 0.2875/(6.023*10^23)

= 4.77*10^-25 kg <<<<<<<------------Answer